OpenStudy (anonymous):

(4x^2+4x-2)+ (-9x^2+ 7)

OpenStudy (anonymous):

ok so the first step is to add like terms

OpenStudy (kc_kennylau):

Well, you can take away the parentheses first

OpenStudy (anonymous):

how would I do that @kc_kennylau

OpenStudy (kc_kennylau):

Just try to write it without the parentheses

OpenStudy (kc_kennylau):

And be careful that (a-b)-(c-d) = a-b-c+d (this case doesn't apply here)

OpenStudy (anonymous):

the problem states that it is addition

OpenStudy (kc_kennylau):

The first step would be like this: (4x^2+4x-2)+ (-9x^2+ 7) =4x^2+4x-2-9x^2+7

OpenStudy (kc_kennylau):

Do you understand this

OpenStudy (anonymous):

combining like terms

OpenStudy (kc_kennylau):

this step is not combining like terms though, but it is the next step

OpenStudy (anonymous):

so I take-1 from the second parentheses?

OpenStudy (kc_kennylau):

Nah

OpenStudy (kc_kennylau):

Or like this: (4x^2+4x-2)+ (-9x^2+ 7) =4x^2+4x-2+ -9x^2+ 7 =4x^2+4x-2-9x^2+7

OpenStudy (anonymous):

confusing me?

OpenStudy (kc_kennylau):

Well, I just took away the parentheses

OpenStudy (anonymous):

k...next step

OpenStudy (kc_kennylau):

Well, we would regroup the terms: 4x^2+4x-2-9x^2+7 =4x^2-9x^2+4x-2+7

OpenStudy (kc_kennylau):

do u understand this

OpenStudy (anonymous):

you distributed -1in the second parentheses

OpenStudy (kc_kennylau):

Well, I don't think I did, since the operation in front of the second parentheses is plus instead of minus

OpenStudy (anonymous):

it looks like you grouped everything (like terms)together...but idk

OpenStudy (kc_kennylau):

well, I just took away the parentheses

OpenStudy (kc_kennylau):

Let's use numbers to understand that: (3+4)+(-9+7) = 3+4-9+7

OpenStudy (anonymous):

ok...you took away the parentheses

OpenStudy (kc_kennylau):

So this is the next step: 4x^2+4x-2-9x^2+7 =4x^2-9x^2+4x-2+7

OpenStudy (kc_kennylau):

Sorry I have to go now